OpenStudy (anonymous):

What is the slope of the line 2x - 5y = 9? Question 11 options: 9 -2/5 -5 2/5

hero (hero):

2x-5y = 9 2x - 9 = 5y (2x-9)/5 = y slope is 2/5

OpenStudy (anonymous):

2/5

OpenStudy (anonymous):

thanks guys